A plot of land has width of x meters. The length of the plot of land is 5 meters more than 3 times it's width. What is the area of the land?
Answer by JBarnum(2044)

A plot of land has width of x meters. The length of the plot of land is 5 meters more than 3 times it's width. What is the area of the land?
x=width L=Length Length x width = area in this case its L times x =area
If x=width and L=5+3x then the equation L+x=A is as follows
(5+3x)(x)=A
3x^2 +5x=A
i cant find enough info to solve the problem further unless the 3x^2+5x is sufficient